OpenStudy (anonymous):

when f(x) is even is the function always 2?

OpenStudy (anonymous):

If f(x) is even, what is \[\int\limits_{-3}^{3}f(x)dx?\] . what is the average value of f(x)on the interval x=-3 to x=3?

OpenStudy (anonymous):

this is the part b) to a question. Part a) told us that \[\int\limits_{0}^{3}f(x)dx=6\]

OpenStudy (mathmate):

If f(x) is an even function, it means that f(x)=f(-x). This means that by symmetry: \(\int_{-b}^b f(x)dx = 2\int_0^b f(x)dx\)

OpenStudy (anonymous):

how did you get that, im kind of lost

OpenStudy (mathmate):

While we're on the subject, an odd function has the property: \(f(-x)=-f(x)\) Example: \(sin(x), tan(x), x^3,x^5\), are all odd, while \(cos(x), sec(x), x^2,x^4,x^6... \)are all even. and hence by symmetry: \int_{-b}^b f(x)dx = 0 

OpenStudy (anonymous):

or is that just a formula?

OpenStudy (mathmate):

It's by symmetry.  As you can see in the drawing, area A = area B, so Area A+B = 2*Area B